# PartSDF: Part-Based Implicit Neural Representation for Composite 3D Shape Parametrization and Optimization

### v1

@@ B.6 Metrics
To evaluate the volume accuracy, we use Intersection over Union (IoU). It is defined as the ratio of volume between the intersection of the shapes over their union. In practice, we compute it using occupancy values on a grid: We find the smallest bounding box that contains the two shapes, then sample points on a $128^{3}$ grid. For each point $\mathbf{x}$, we get the binary occupancy of each shape as $o_{\mathcal{S}_{1}}(\mathbf{x})$ and $o_{\mathcal{S}_{2}}(\mathbf{x})$, where $o_{\mathcal{S}}(\mathbf{x})$ is $1$ if $\mathbf{x}$ is inside the shape and 00 otherwise. The IoU is then computed as
@@
Finally, we use Image Consistency (IC) (Guillard et al., 2022) to evaluate the appearance and surface normals, using renderings of the shapes. Let $\mathcal{K}$ be a set of 8 cameras located at the vertices of a cuboid that encompasses the shape, looking at its centroid. For each camera $k\in\mathcal{K}$, we render the binary silhouettes $S_{k}\in\{0,1\}^{256\times 256}$ and normal map $N_{k}\in^{256\times 256\times 3}$ of shape $\mathcal{S}_{1}$, and similarly $\tilde{S}_{k}$ and $\tilde{N}_{k}$ for shape $\mathcal{S}_{2}$. The IC between these shapes is defined as
@@
We additionally report per-part average IoU (pIoU) as a way to measure part reconstruction accuracy, noting that CD and IC might have issues if the ground-truth parts are not watertight. In order to compute the occupancy of an open part, we rely on a strategy similar to our SDF supervision, see Figure 3. First, the occupancy of the full shape is computed on the grid. Then, for each point where the occupancy is $1$, we look for the closest part to it. This part will get an occupancy of $1$ at that point and all the others will get 00. Therefore, part occupancy can be seen as the intersection of the full shape’s occupancy with the part’s ”closest region” as visualized in Figure 3(b). Note that for baselines without part correspondences, i.e. DAE-Net and PQ-Net, we must first match the reconstructed parts to the ground truth (GT) ones: For each shape and each reconstructed part, we match it with the GT part with which it has the highest IoU.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix C Part Optimization
As described in Section 4.4, we optimize shapes to demonstrate how PartSDF is used as a part-aware shape prior. In this section, we provide further experimental details and results.
@@
We use a graph convolutional neural network (GCNN), specifically with GraphSage layers (Hamilton et al., 2017), to predict the surface pressure distribution of a car. The predicted surface pressure is integrated to compute the drag force (Equation 9), which is then normalized by the mass density of the fluid, its velocity squared, and the frontal surface area of the shape to obtain $C_{d}$. We simulate the cars in our dataset, following the well-established setup of the DrivAer model (Wieser et al., 2014), to compute the corresponding surface pressures and $C_{d}$ values. The dataset thus created is used to trained the GCNN-based surrogate model.
@@
The optimization minimizes the integrated pressure drag predicted by the surrogate model, backpropagating through the meshing to adjust the part parameters (Equation 10). Regularization terms $\mathcal{L}_{\text{reg}}$ are applied to maintain consistency in the latent space and prevent excessive deviations:
@@ Results.
With this experimental setup, we optimize multiple cars from our dataset and simulate the resulting shapes in OpenFOAM (OpenCFD, ) to obtain their final drag coefficient $C_{d}$. Because simulations are expensive ($>10$h depending on the meshing resolution of the 3D space), we must limit ourselves to a subset of 35 cars. The average drag coefficients before and after optimizations are
@@
with an average relative improvement of 12.7%. We visualize some optimization results in Figure 9. As can be observed, only the body of the car has been modified while the wheels are unchanged. Nonetheless, they stay consistent with each other and no inter-penetration happens.
@@ Appendix D Part Interpolation
To demonstrate the smoothness and consistency of our latent and pose space, we perform an interpolation experiment. Given pair of shapes, we linearly interpolate between their part latent vectors and poses, with the exception of rotation quaternions for which we use spherical linear interpolation (slerp). At multiple interpolation steps, we reconstruct the intermediate shapes using our decoder and visualize such interpolation ”path” in Figure 10, illustrating smooth transitions between the pair of shapes. Notably, the part-based structure remains coherent throughout the interpolation. This highlights the effectiveness of our learned latent and pose spaces in representing composite shapes.
Appendix E Ablation Study
Table 3: Ablation on PartSDF components. We ablate the use of part poses, latent modulation (latMod) and convolutional layers (conv) on the Car dataset. Best performance is achieved with poses and latent modulation, though we note that convolutions decrease slightly the part accuracy of our model. However, this is counterbalanced by enabling part manipulation, as in Figure 11.
CD ($\downarrow$) IoU (%, $\uparrow$) IC ($\uparrow$) pIoU (%, $\uparrow$)
w/o poses 1.43 97.81 0.928 91.71
w/o latMod 1.39 97.37 0.923 92.33
w/o conv 1.27 98.02 0.932 96.37
PartSDF 1.27 98.02 0.931 94.89
Refer to caption Refer to caption
w/o $h_{\text{fc}}$ w/ $h_{\text{fc}}$
Refer to caption Refer to caption Refer to caption
GT w/o $\mathcal{L}_{\text{inter}}$ w/ $\mathcal{L}_{\text{inter}}$
(a) Convolution layers (b) Non-intersection loss
Figure 11: Ablation study of PartSDF. (a) Without the convolution layers $h_{\text{fc}}$ (Section 3.1), manipulating a part does not affect the others: The translated and scaled wheels incorrectly collide with the car. With these added layers, manipulating some parts correctly affect the others: The car adapts to the new wheel poses without intersection. (b) Without the non-intersection loss $\mathcal{L}_{\text{inter}}$, the parts incorrectly overlaps. With it, the model learns the correct part decomposition.
We conduct an ablation study to evaluate the impact of convolutional layers, the use of pose parameters, and latent modulation in PartSDF, and report metrics on the Car dataset in Table 3. Pose parameters allow independent control of part transformations, which improves the shape reconstruction performance but also allows more precise part manipulation, while the addition of latent modulation is a simple modification that improves surface and part reconstruction, as measured by Chamfer-Distance and part IoU. On the other hand, convolutional layers cause a slight decrease in part accuracy, largely counterbalanced by the manipulation capability it enables. In Figure 11, we show that adding these layers to PartSDF maintains the consistency between parts when doing part manipulation. Additionally, we show in Figure 11 the necessity of the non-intersection loss $\mathcal{L}_{\text{inter}}$: without it, the reconstruction losses do not prevent the parts from bleeding into one another, causing the model to learn an incorrect part decomposition with overlapping parts.
Appendix F Additional Experimental Results
In this section, we provide additional results and visualization on shape reconstruction (Section F.1), generation (Section F.2) and manipulation (Section F.3).
F.1 Shape Reconstruction
We provide additional examples of shape reconstruction on the test set in Figure 12. We also visualize the part poses as primitives and results using the point encoder. DAE-Net and BAE-Net struggle to recover thin parts and complex shapes, while PQ-Net tend to inflates them without details. On the opposite, PartSDF recovers accurately the shapes and the parts. We note that with a point cloud encoder, the results may not be as accurate, e.g., the last chair where the armrest segmentation is different, though it is still sensible.
3DShape2VecSet Ours-1P DAE-Net BAE-Net PQ-Net Ours-Prims Ours Ours-PC GT
@@ F.1 Shape Reconstruction
Figure 12: Additional shape reconstruction. Reconstructions of test shapes using either the baselines or our model. We use a single part with Ours-1P, we visualize the pose parts of Ours as cuboid or cylinders in Ours-Prims, and Ours-PC is the reconstruction using the point cloud encoder, as described in Section 4.1. For part-based method, we translate the helix outside of the mixers for visualization.
F.2 Shape Generation
In Figure 13, we visualize generation examples with PQ-Net and PartSDF. We also show the part latent generation conditioned on poses for our model. The generated shapes with our method in combination of SALAD (Koo et al., 2023) are more detailed than PQ-Net + a latentGAN, with greater variety for cars and mixers (as reported in Table 2).
@@ F.2 Shape Generation
Figure 13: Additional shape generation. Randomly generated shapes and their parts on all datasets. We also provide examples of pose conditioned generation (Ours) where part latents are generated based on the poses’ coarse description of the shape (top image for each pair). The helix is translated outside of the mixers for visualization.
F.3 Shape Manipulation
Finally, we perform more manipulation in Figure 14 with PartSDF on all dataset. By changing part latents, the affected parts have modified appearances but adapt to the original pose and to each others. For example, the car bodies have the same general scales but their wheel wells fit correctly the original wheels, the mixers helices fit the original tube, and the chairs keep the same overall structure while each parts are locally different. When editing their poses, the parts are moved and deformed, but also preserve their original features, e.g., the cars keep their original shapes but adapt to their new wheels, the mixers and their helices are rescaled but conserve their inside/outside relationship, and the size and height of chairs can be changed while maintaining the chairs’ features. Note that this requires the part poses to be correct and coherent. Indeed, if they are manually set to be out of contact and disorganized in 3D space, the resulting shape will be meaningless.

### v2

@@ B.6 Metrics
To evaluate the volume accuracy, we use Intersection over Union (IoU). It is defined as the ratio of volume between the intersection of the shapes over their union. In practice, we compute it using occupancy values on a grid: We find the smallest bounding box that contains the two shapes, then sample points on a $128^{3}$ grid. For each point $\mathbf{x}$, we get the binary occupancy of each shape as $o_{\mathcal{S}_{1}}(\mathbf{x})$ and $o_{\mathcal{S}_{2}}(\mathbf{x})$, where $o_{\mathcal{S}}(\mathbf{x})$ is $1$ if $\mathbf{x}$ is inside the shape and 0 otherwise. The IoU is then computed as
@@
Finally, we use Image Consistency (IC) (Guillard et al., 2022) to evaluate the appearance and surface normals, using renderings of the shapes. Let $\mathcal{K}$ be a set of 8 cameras located at the vertices of a cuboid that encompasses the shape, looking at its centroid. For each camera $k\in\mathcal{K}$, we render the binary silhouettes $S_{k}\in\{0,1\}^{256\times 256}$ and normal map $N_{k}\in\mathbb{R}^{256\times 256\times 3}$ of shape $\mathcal{S}_{1}$, and similarly $\tilde{S}_{k}$ and $\tilde{N}_{k}$ for shape $\mathcal{S}_{2}$. The IC between these shapes is defined as
@@
We additionally report per-part average IoU (pIoU) as a way to measure part reconstruction accuracy, noting that CD and IC might have issues if the ground-truth parts are not watertight. In order to compute the occupancy of an open part, we rely on a strategy similar to our SDF supervision, see Figure˜4. First, the occupancy of the full shape is computed on the grid. Then, for each point where the occupancy is $1$, we look for the closest part to it. This part will get an occupancy of $1$ at that point and all the others will get 0. Therefore, part occupancy can be seen as the intersection of the full shape’s occupancy with the part’s "closest region" as visualized in Figure˜4(b). Note that for baselines without part correspondences, i.e. DAE-Net and PQ-Net, we must first match the reconstructed parts to the ground truth (GT) ones: For each shape and each reconstructed part, we match it with the GT part with which it has the highest IoU.
@@ Appendix C Part Optimization
As described in Section˜4.4, we optimize shapes to demonstrate how PartSDF is used as a part-aware shape prior. In this section, we provide further experimental details and results.
@@
We use a graph convolutional neural network (GCNN), specifically with GraphSage layers (Hamilton et al., 2017), to predict the surface pressure distribution of a car. The predicted surface pressure is integrated to compute the drag force (Equation˜10), which is then normalized by the mass density of the fluid, its velocity squared, and the frontal surface area of the shape to obtain $C_{d}$. We simulate the cars in our dataset, following the well-established setup of the DrivAer model (Wieser et al., 2014), to compute the corresponding surface pressures and $C_{d}$ values. The dataset thus created is used to trained the GCNN-based surrogate model.
@@
The optimization minimizes the integrated pressure drag predicted by the surrogate model, backpropagating through the meshing to adjust the part parameters (Equation˜11). Regularization terms $\mathcal{L}_{\text{reg}}$ are applied to maintain consistency in the latent space and prevent excessive deviations:
@@ Results.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
With this experimental setup, we optimize multiple cars from our dataset and simulate the resulting shapes in OpenFOAM (OpenCFD, ) to obtain their final drag coefficient $C_{d}$. Because simulations are expensive (>10>10h depending on the meshing resolution of the 3D space), we must limit ourselves to a subset of 35 cars. The average drag coefficients before and after optimizations are
@@
with an average relative improvement of 12.7%. We visualize some optimization results in Figure˜10. As can be observed, only the body of the car has been modified while the wheels are unchanged. Nonetheless, they stay consistent with each other and no inter-penetration happens.
@@ Appendix D Part Interpolation
To demonstrate the smoothness and consistency of our latent and pose space, we perform an interpolation experiment. Given pair of shapes, we linearly interpolate between their part latent vectors and poses, with the exception of rotation quaternions for which we use spherical linear interpolation (slerp). At multiple interpolation steps, we reconstruct the intermediate shapes using our decoder and visualize such interpolation "path" in Figure˜11, illustrating smooth transitions between the pair of shapes. Notably, the part-based structure remains coherent throughout the interpolation. This highlights the effectiveness of our learned latent and pose spaces in representing composite shapes.
Appendix E Additional Experimental Results
In this section, we provide additional results and visualization on shape reconstruction (Section˜E.1), generation (Section˜E.2) and manipulation (Section˜E.3).
E.1 Shape Reconstruction
We provide additional examples of shape reconstruction on the test set in Figure˜12. We also visualize the part poses as primitives and results using the point encoder. DAE-Net and BAE-Net struggle to recover thin parts and complex shapes, while PQ-Net tend to inflates them without details. On the opposite, PartSDF recovers accurately the shapes and the parts. We note that with a point cloud encoder, the results may not be as accurate, e.g., the last chair where the armrest segmentation is different, though it is still sensible.
3DShape2VecSet Ours-1P DAE-Net BAE-Net PQ-Net PASTA Ours Ours-PC GT
@@ E.1 Shape Reconstruction
Figure 12: Additional shape reconstruction. Reconstructions of test shapes using the baselines and our model. We use a single part with Ours-1P and Ours-PC is the reconstruction using the point cloud encoder, as described in Section˜4.1. For part-based method, we translate the helix outside of the mixers for visualization.
E.2 Shape Generation
In Figure˜13, we visualize generation examples with PQ-Net and PartSDF. We also show the part latent generation conditioned on poses for our model. The generated shapes with our method in combination of SALAD (Koo et al., 2023) are more detailed than PQ-Net + a latentGAN, with greater variety for cars and mixers (as reported in Table˜2).
@@ E.2 Shape Generation
Figure 13: Additional shape generation. Randomly generated shapes and their parts on all datasets. We also provide examples of pose conditioned generation (Ours) where part latents are generated based on the poses’ coarse description of the shape (top image for each pair). When possible, the helix is translated outside of the mixers for visualization.
E.3 Shape Manipulation
Finally, we perform more manipulation in Figure˜14 with PartSDF on all dataset. By changing part latents, the affected parts have modified appearances but adapt to the original pose and to each others. For example, the car bodies have the same general scales but their wheel wells fit correctly the original wheels, the mixers helices fit the original tube, and the chairs keep the same overall structure while each parts are locally different. When editing their poses, the parts are moved and deformed, but also preserve their original features, e.g., the cars keep their original shapes but adapt to their new wheels, the mixers and their helices are rescaled but conserve their inside/outside relationship, and the size and height of chairs can be changed while maintaining the chairs’ features. Note that this requires the part poses to be correct and coherent. Indeed, if they are manually set to be out of contact and disorganized in 3D space, the resulting shape will be meaningless.